OpenStudy (anonymous):

can someone explain the geometric implications of taking the exponential of a matrix please?

OpenStudy (anonymous):

it was introduced to me in a problem a few days ago

OpenStudy (kinggeorge):

Suppose your matrix translates to some sort of linear action on R^2 (or similar). The exponential is the same thing as applying the transformation twice.

OpenStudy (kinggeorge):

or thrice. Or four times.

OpenStudy (anonymous):

sorry i mean exp(A) not A^n

OpenStudy (kinggeorge):

nevermind :(

OpenStudy (anonymous):

\[\exp(A) =\sum{\frac{1}{r!} A^r}\]

OpenStudy (anonymous):

how do i type infinity in the equation thing?

OpenStudy (kinggeorge):

\infty

OpenStudy (amistre64):

.... how odd if i were to take a blind stab at this: e^A = B A = ln(B) .... and thats as far as I get on that idea lol

OpenStudy (zarkon):

just using the trylor expansion for e

OpenStudy (zarkon):

*\(e^x\)

OpenStudy (anonymous):

i'll find the question, i did it fine, i just want extra explanation

OpenStudy (amistre64):

A^r = P-1 D P right? if its diag.able

OpenStudy (amistre64):

D^r that is

OpenStudy (anonymous):

sorry, im answering a question i wont be long :)

OpenStudy (anonymous):

i solved it, i was just interested in getting more info on exponentials of matrices as this is the first i have encountered

OpenStudy (anonymous):

Square matrices?

OpenStudy (kinggeorge):

It seems like you might want to consider your matrix \(M\) as an equivalent to \(i=\sqrt{-1}\) and from there draw some parallels to the complex plane.

OpenStudy (anonymous):

yes. i am also interested in what happens if its a general matrix M instead of that specific matrix

OpenStudy (kinggeorge):

I honestly don't know how you could visualize exp(A) with a generalized matrix matrix of size nxn. Unless \(n=1\). Then it's just plain old \(e^a\) for some \(a\).

OpenStudy (anonymous):

The matrix exponential for square matrices has a lot of relevance in rotation theory..

OpenStudy (anonymous):

http://en.wikipedia.org/wiki/Matrix_exponential

OpenStudy (anonymous):

this maths is so much more exciting than school work!